# Feedback-driven quantum reservoir computing for time-series analysis

### v1

@@ Abstract
Quantum reservoir computing (QRC) is a highly promising computational paradigm that leverages quantum systems as a computational resource for nonlinear information processing. While its application to time-series analysis is eagerly anticipated, prevailing approaches suffer from the collapse of the quantum state upon measurement, resulting in the erasure of temporal input memories. Neither repeated initializations nor weak measurements offer a fundamental solution, as the former escalates time complexity while the latter restricts the information extraction from the Hilbert space. To address this issue, we propose a feedback-driven QRC framework. This methodology employs projective measurements on all qubits for unrestricted access to the quantum state, with the measurement outcomes subsequently fed back into the reservoir to restore the memory of prior inputs. We demonstrate that our QRC successfully acquires the fading-memory property through the feedback, a critical element in time-series processing. Notably, analysis of measurement trajectories reveal three distinct phases depending on the feedback strength, with the memory performance maximized at the edge of chaos. We also evaluate the predictive capabilities of our QRC, demonstrating its suitability for forecasting signals originating from quantum spin systems.
@@ I INTRODUCTION
In the modern information society, quantum machine learning is gaining heightened interest for its potential in advanced data processing [1, 2, 3, 4, 5, 6, 7, 8]. The cornerstone of the advantage lies in the use of the exponentially large Hilbert space, which allows for the accommodation of high-dimensional features of datasets and the emulation of a substantial number of computational neurons [9, 10, 11, 12]. A quantum extension of the reservoir computing paradigm, known as quantum reservoir computing (QRC), also leverages this high dimensionality [13, 14]. Central to the philosophy of (general) reservoir computing is its distinctive architecture, where a fixed, randomly connected network of nodes–“reservoir”–processes input data by projecting it onto an internal feature space [15, 16, 17, 18, 19, 20, 21]. This strategy limits the training parameters only to the weights on the read-out part of the network, significantly reducing optimization costs and enabling high-speed, real-time temporal processing. Previous investigations have demonstrated that computational capabilities tend to be maximized when the reservoir part resides proximate to the boundary separating a stable and an unstable dynamical regimes, commonly referred to as the edge of chaos [22, 23, 24, 25, 26, 27]. Importantly, given the fixed nature of the internal parameters, the reservoir part can be realized either with an artificial recurrent neural network or with physical systems. In the QRC framework, a quantum system is harnessed as the physical embodiment of the reservoir part, termed a quantum reservoir. The intrinsic nonlinearity and high dimensionality of quantum systems perfectly align with the heuristic requirements for effective reservoir computing systems, which facilitate the extraction and transformation of intricate data features [13, 14, 19]. Thus far, theoretical proposals of the QRC have yielded promising results in addressing temporal and classification tasks [28, 29, 30, 31, 32, 33], and experimental demonstrations have been reported on several quantum reservoir settings, including nuclear magnetic resonance systems [34] and superconducting qubits [35]. Furthermore, since the information processing capabilities of the QRC are fundamentally linked to the nature of the quantum reservoirs [36, 37], it can be employed as a tool to investigate the quantum system itself, a concept recently formalized as quantum reservoir probing [38, 39].
@@
Particularly focusing on temporal information processing, the fading-memory property emerges as a crucial aspect of the reservoir computing framework. This property dictates that when the reservoir projects input data onto the feature space, the temporal dependencies between different inputs are faithfully mapped within that internal space, and the influences of earlier inputs, being less significant than more recent ones, gradually diminish [15, 16]. Herein lies a significant challenge for the QRC: the inherent conflict between information extraction and memory preservation within the quantum reservoir. Conventional QRC architectures rely on quantum measurements to extract information from the quantum reservoir. However, these measurements inevitably induce measurement back-action, altering the quantum state of the reservoir. As a result, the memories of previous inputs encoded within the quantum reservoir are completely erased. This disturbance compromises the preservation of short-term memory and thus hinders the ability to process temporal information due to the absence of the crucial fading-memory property.
@@
Figure 1: Architectures of QRC with an input sequence $\{s_{k}\}$ and an initial state $|\psi_{0}\rangle$. $\mathcal{L}$ symbolizes a generalized operation for a single input cycle parameterized by external inputs, the specific form of which depends on each realization. (a) Quadratic restarting protocol where the entire input signal is provided at every cycles to reinitialize the reservoir after the measurements. (b) Linear restarting protocol where the last $\tau$ inputs are repeated, assuming less effects from further previous inputs. (c) Online protocol that continuously monitors the time evolution using weak measurements or projective measurements only on limitied part of the system. (d) Feedback protocol where the measurement result at the $(k-1)$-th cycle determines the operation $\mathcal{L}$ at the $k$-th cycle.
@@
To address this monitoring issue, a common strategy involves repeated initialization after each measurement. The entire process is executed from the beginning for every individual input, thereby effectively neglecting the effects of measurement back-actions [Fig. 1(a)] [35, 40, 41, 42]. However, this approach naturally necessitates $O(L^{2})$ input operations to address a time-series of length $L$, leading to a computationally expensive quadratic growth in execution time. A potential reduction in temporal complexity could be achieved under the premise of the fading-memory property. As the reservoir gradually loses information about distant past inputs, the simplification of the process by limiting input repetitions to recent cycles is expected to have minimal impact on memory performance [Fig. 1(b)]. By initiating the dynamics for the input cycle $k$ from the cycle $k-\tau+1$ rather than the cycle $1$, the total number of input repetitions reduces to $\tau L$, yielding a time complexity of $O(L)$ [43]. Nevertheless, despite this improvement, the requirement for repeating gate operations to input identical data still significantly compromises the real-time processing capabilities, a key advantage of the reservoir computing paradigm.
@@
A promising approach seeks to balance short-term memory preservation with real-time functionality, by continuously monitoring quantum dynamics online while partially retaining the reservoir’s quantum state at each measurement [Fig. 1(c)]. By employing weak measurements or projective measurements on a limited subset of qubits, the QRC generates output online while mitigating the erasure of past input memory [44, 45, 46, 47, 48]. However, these approaches only permit limited information extraction, as partial measurements restrict comprehensive access to the quantum state of the reservoir. This constraint hinders the full utilization of the exponentially large Hilbert space in which the quantum reservoir resides, thus undermining a fundamental advantage of the QRC frameworks. Additionally, the necessity for supplementary ancillary systems in weak measurement configurations increases overall system complexity. Furthermore, considering the coherence time, there would be an upper bound on the length $L$ of the input data for which the system can maintain online. Consequently, the development of a simple and versatile online QRC protocol that fully leverages the quantum reservoir systems remains an urgent area of research.
@@
In this paper, we propose a feedback-driven online QRC protocol that achieves complete access to the quantum state through projective measurements on all qubits [Fig. 1(d)]. While these measurements destroy short-term memory within the reservoir’s state, information of prior inputs is effectively revived by feedback connections that incorporate the measurement outcomes back into the quantum reservoir [49, 50, 51, 52]. By evaluating the short-term memory performance, we demonstrate that our QRC framework successfully acquire the fading-memory property through the feedback, showcasing its applicability to time-series analysis. Notably, the proposed QRC system is highly tunable, as the strength of input and feedback, which strongly relates to information embedding and memory preservation, can be controlled externally. Indeed, analysis of the measurement trajectories reveals three distinct phases depending on the feedback strength. The memory performance peaks at the phase boundary between the stable and the unstable phase, i.e., the edge of chaos. We further evaluate the predictive performance of our QRC for various time-series signals. Benchmark analysis against a conventional classical reservoir computing framework reveals that our QRC is highly suitable for forecasting signals derived from quantum spin systems. Our findings highlight the substantial potential of the QRC in the burgeoning quantum era as a potent instrument for time-series analysis.
@@ II.1 Architecture of the feedback-driven QRC
In step (i) of the $k$-th cycle, the input value $s_{k}$ is injected into the reservoir via the application of a two-qubit gate $\mathcal{R}_{1,2}(a_{\mathrm{in}}s_{k})$ on qubits $1$ and $2$, where $a_{\mathrm{in}}$ denote an input scaling weight. The operator $\mathcal{R}_{i,j}$ is defined as
@@
For clarity, we next describe the step (iv) of the ($k-1$)-th cycle. Here, Pauli $Z$ measurements are executed on all qubits, which yield the measurement result vector $\bm{z}_{k-1}=[\langle Z_{1}\rangle_{k-1},\langle Z_{2}\rangle_{k-1},\dots,\langle Z_{n}\rangle_{k-1}]^{\top}$ (where $Z_{i}$ denotes the Pauli $Z$ for the qubit $i$). Returning to step (ii) of the $k$-th cycle, the vector $\bm{z}_{k-1}$ is provided back into the reservoir through the feedback connection. The operator $\mathcal{R}_{i,j}(a_{\mathrm{fb}}z_{k-1}^{\alpha})$ is employed for feedback with weight $a_{\mathrm{fb}}$, which encodes all components of $\bm{z}_{k-1}$ according to the configuration shown in Fig. 2(a). Although the projective measurement strongly perturbs the quantum state, the feedback connections operate to systematically restore the memory of the provided input history.
@@
Subsequent to the input and feedback processes, a fixed eight-qubit gate $U_{\mathrm{res}}$ is applied in step (iii) to generate entanglement among all qubits. This gate serves as the reservoir component of our QRC architecture, which simultaneously transforms the input and feedback information in a non-linear manner. We employ a Haar random unitary matrix for $U_{\mathrm{res}}$ in the following calculations. Finally, step (iv) concludes the $k$-th cycle with projective measurements on all qubits. The measurement result $\bm{z}_{k}$ is further cycled back to the reservoir via the feedback gate for processing the subsequent $(k+1)$-th input.
@@
Figure 2: (a) Circuit diagram of the feedback-driven QRC architecture at the $k$-th cycle. The red and blue background color represent the gates for the input [step (i)] and feedback [step (ii)] parts, respectively. The input value $s_{k}$ is provided to qubits $1$ and $2$ via $\mathcal{R}_{1,2}(a_{\mathrm{in}}s_{k})$ with the input weight $a_{\mathrm{in}}$. The feedback of the measurement result $\bm{z}_{k-1}$ is given to qubits $3$ to $8$ via $\mathcal{R}_{i,j}(a_{\mathrm{fb}}z_{k-1}^{\alpha})$. Subsequently, the entanglement among all qubits is introduced by the reservoir gate $U_{\mathrm{res}}$ [step (iii)], and all qubits are then measured in the projective manner at the end of the circuit [step (iv)]. (b) Circuit representation of the two-qubit operator $\mathcal{R}$ defined in Eq. (1).
@@
It is worth highlighting a key advantage of our architecture in terms of time complexity, which arises from the intrinsic independence of quantum circuits over different cycles. In our QRC, the quantum state at the $k$-th cycle depends exclusively on the current input value $s_{k}$ and the preceding measurement outcome $\bm{z}_{k-1}$, explicitly represented as $|\psi(s_{k},\bm{z}_{k-1})\rangle$. Even after the measurement, this state can be efficiently reconstructed by applying a limited set of gates, leveraging $\bm{z}_{k-1}$ stored in external memory; there is no need to re-execute the $(k-1)$-th and prior cycles when addressing the $k$-th input. Consequently, our protocol offers remarkable operational flexibility, enabling features such as starting from an intermediate cycle or running different cycles on different devices. In addition, owing to the frequent measurement and subsequent initialization, a coherence time sufficient for the execution of a single cycle is adequate, regardless of the length of the input data. These aspects stand in stark contrast to other implementations depicted in Figs. 1(a)-(c), which, due to their dependence on the input sequence as $|\psi(s_{k},s_{k-1},s_{k-2}\dots)\rangle$, necessitate significantly longer, uninterrupted operations.
@@ II.2 Learning procedure
In the QRC protocol, the final output $y_{k}$ at the $k$-th cycle is obtained through a linear transformation of the measurement result $\bm{z}_{k}$. Here, a specific task for the QRC is formulated by setting a target output $\bar{y}_{k}$ based on the input sequence $\{s_{k}\}$. Our objective in the output part is to construct a regression model from the collection of measurement outcomes that accurately approximates the target output over various cycles. As a prerequisite, we define $\bm{z}_{0}$, required for the first cycle ($k=1$), as a random vector with each element sampled uniformly from the interval $[0,1]$. The influence of this initial condition is disregarded by discarding the initial $l_{\mathrm{w}}$ cycles.
@@
Following the training phase, we construct an $[l_{\mathrm{ts}}\times(n+1)]$-dimensional matrix ${X}_{\mathrm{ts}}$ for testing in a manner analogous to Eq. (2) using the measurement outcomes from the ($l_{\mathrm{w}}+l_{\mathrm{tr}}+1$)-th cycle to the ($l_{\mathrm{w}}+l_{\mathrm{tr}}+l_{\mathrm{ts}}$)-th cycle. The output is then calculated as $\bm{y}_{\mathrm{ts}}={X}_{\mathrm{ts}}\bm{w}$ using the trained weight. The similarity between the output $\bm{y}_{\mathrm{ts}}$ and the target output $\bar{\bm{y}}_{\mathrm{ts}}$ in the testing phase signifies the reservoir performance. As a quantitative metric, we employ the determination coefficient
$R^{2}=\frac{\mathrm{cov}(\bar{\bm{y}}_{\mathrm{ts}},\bm{y}_{\mathrm{ts}})}{\sigma^{2}(\bar{\bm{y}}_{\mathrm{ts}})\sigma^{2}(\bm{y}_{\mathrm{ts}})},$ (4)

### v2

@@ Abstract
Quantum reservoir computing (QRC) is a highly promising computational paradigm that leverages quantum systems as a computational resource for nonlinear information processing. While its application to time-series analysis is eagerly anticipated, prevailing approaches suffer from the collapse of the quantum state upon measurement, resulting in the erasure of temporal input memories. Neither repeated initializations nor weak measurements offer a fundamental solution, as the former escalates the time complexity while the latter restricts the information extraction from the Hilbert space. To address this issue, we propose the feedback-driven QRC framework. This methodology employs projective measurements on all qubits for unrestricted access to the quantum state, with the measurement outcomes subsequently fed back into the reservoir to restore the memory of prior inputs. We demonstrate that our QRC successfully acquires the fading-memory property through the feedback connections, a critical element in time-series processing. Notably, analysis of measurement trajectories reveal three distinct phases depending on the feedback strength, with the memory performance maximized at the edge of chaos. We also evaluate the predictive capabilities of our QRC, demonstrating its suitability for forecasting signals originating from quantum spin systems.
@@ I INTRODUCTION
In the modern information society, quantum machine learning is gaining heightened interest for its potential in advanced data processing [1, 2, 3, 4, 5, 6, 7, 8, 9]. The cornerstone of the advantage lies in the use of the exponentially large Hilbert space, which allows for the accommodation of high-dimensional features of datasets and the emulation of a substantial number of computational neurons [10, 11, 12, 13]. A quantum extension of the reservoir computing paradigm, known as quantum reservoir computing (QRC), also leverages this high dimensionality [14, 15]. Central to the philosophy of (general) reservoir computing is its distinctive architecture, where a fixed, randomly connected network of nodes–“reservoir”–processes input data by projecting it onto an internal feature space [16, 17, 18, 19, 20, 21, 22]. This strategy limits the training parameters only to the weights on the read-out part of the network, significantly reducing optimization costs and enabling high-speed, real-time temporal processing. Previous investigations have demonstrated that computational capabilities tend to be maximized when the reservoir part resides proximate to the boundary separating a stable and an unstable dynamical regimes, commonly referred to as the edge of chaos [23, 24, 25, 26, 27, 28]. Importantly, given the fixed nature of the internal parameters, the reservoir part can be realized either with artificial recurrent neural networks or with physical systems. In the QRC framework, a quantum system is harnessed as the physical embodiment of the reservoir part, termed a quantum reservoir. The intrinsic nonlinearity and high dimensionality of quantum systems perfectly align with the heuristic requirements for effective reservoir computing systems, which facilitate the extraction and transformation of intricate data features [14, 15, 20]. Thus far, theoretical proposals of the QRC have yielded promising results in addressing temporal and classification tasks [29, 30, 31, 32, 33, 34, 35, 36], and experimental demonstrations have been reported on several quantum reservoir settings, including nuclear magnetic resonance (NMR) systems [37] and superconducting qubits [38]. Furthermore, since the information processing capabilities of the QRC are fundamentally linked to the nature of the quantum reservoirs [39, 40], it can be employed as a tool to investigate the quantum system itself, a concept recently formalized as quantum reservoir probing [41, 42].
@@
Particularly focusing on temporal information processing, the fading-memory property emerges as a crucial aspect of the reservoir computing framework. This property dictates that when the reservoir projects input data onto the feature space, the temporal dependencies between different inputs are faithfully mapped within that internal space, and the influences of earlier inputs, being less significant than more recent ones, gradually diminish [16, 17]. Herein lies a significant challenge for the QRC: the inherent conflict between information extraction and memory preservation within the quantum reservoir. Conventional QRC architectures rely on quantum measurements to extract information from the quantum reservoir. However, these measurements inevitably induce measurement back-action, leading to the destruction of the quantum state of the reservoir. As a result, the memories of previous inputs encoded within the quantum reservoir are completely erased. This disturbance compromises the preservation of short-term memory and thus hinders the ability to process temporal information due to the absence of the crucial fading-memory property.
@@
Figure 1: Architectures of QRC with an input sequence $\{s_{k}\}$ and an initial state $|\psi_{0}\rangle$. $\mathcal{L}$ symbolizes a generalized operation for a single input cycle parameterized by external inputs, the specific form of which depends on each realization. (a) Quadratic restarting protocol where the entire input signal is provided at every cycles to reinitialize the reservoir after the measurements. (b) Linear restarting protocol where the last $\tau$ inputs are repeated, assuming less effects from further previous inputs. (c) Online protocol that continuously monitors the time evolution using weak measurements or projective measurements only on limited part of the system. (d) Feedback protocol where the measurement result at the $(k-1)$-th cycle determines the operation $\mathcal{L}$ at the $k$-th cycle.
@@
To address this monitoring issue, a common strategy involves repeated initialization after each measurement. The entire process is executed from the beginning for every individual input, thereby effectively neglecting the effects of measurement back-actions [Fig. 1(a)] [38, 43, 44, 45]. However, this approach naturally necessitates $O(L^{2})$ input operations to address a time-series of length $L$, leading to a computationally expensive quadratic growth in execution time. A potential reduction in temporal complexity could be achieved under the premise of the fading-memory property. As the reservoir gradually loses information about distant past inputs, the simplification of the process by limiting input repetitions to recent cycles is expected to have minimal impact on reservoir performance [Fig. 1(b)]. By initiating the dynamics for the input cycle $k$ from the cycle $k-\tau+1$ rather than the cycle $1$, the total number of input repetitions reduces to $\tau L$, yielding a time complexity of $O(L)$ [46]. Nevertheless, despite this improvement, the requirement for repeating gate operations to input identical data still significantly compromises the real-time processing capabilities, a key advantage of the reservoir computing paradigm.
@@
A promising approach seeks to balance short-term memory preservation with real-time functionality, by continuously monitoring quantum dynamics online while partially retaining the reservoir’s quantum state at each measurement [Fig. 1(c)]. By employing weak measurements or projective measurements on a limited subset of qubits, the QRC generates output online while mitigating the erasure of past input memory [47, 48, 49]. Continuous variable quantum optical systems offer efficient platforms for this architecture [50, 51, 52]. However, these approaches only permit limited information extraction, as partial measurements do not provide comprehensive access to the quantum state of the reservoir. This constraint hinders the full utilization of the Hilbert space in which the quantum reservoir resides. Given that the exponential scaling of dimensionality with system size is a fundamental strength of the QRC, this may constitute a drawback when compared to alternative configurations of comparable system size, although, the extent to which it negatively affects performance would depend on the specific implementations. Consequently, the development of a simple and versatile online QRC protocol that fully leverages the Hilbert space of the quantum reservoir remains an urgent area of research.
@@
In this paper, we propose a feedback-driven online QRC protocol that achieves complete access to the quantum state through projective measurements on all qubits [Fig. 1(d)]. While these measurements destroy short-term memory within the reservoir’s state, information of prior inputs is effectively revived by feedback connections that incorporate the measurement outcomes back into the quantum reservoir [53, 54, 55]. By evaluating the short-term memory performance, we demonstrate that our QRC framework successfully acquire the fading-memory property through the feedback, showcasing its applicability to time-series analysis. The proposed QRC system is highly tunable, as the strength of input and feedback, which strongly relates to information embedding and memory preservation, can be controlled externally. Notably, analysis of the measurement trajectories reveals three distinct phases depending on the feedback strength. The memory performance peaks at the phase boundary between the stable and the unstable phase, i.e., the edge of chaos. We further evaluate the predictive performance of our QRC for various time-series signals. Benchmark analysis against a conventional classical reservoir computing framework reveals that our QRC is highly suitable for forecasting signals derived from quantum spin systems. Our findings highlight the substantial potential of the QRC in the burgeoning quantum era as a potent instrument for time-series analysis.
@@ II.1 Architecture of the feedback-driven QRC
In step (i) of the $k$-th cycle, the input value $s_{k}$ is injected into the reservoir via the application of a two-qubit gate $\mathcal{R}_{1,2}(a_{\mathrm{in}}s_{k})$ on qubits $1$ and $2$, where $a_{\mathrm{in}}$ denotes an input scaling weight. The operator $\mathcal{R}_{i,j}$ is defined as
@@
For clarity, we next describe the step (iv) of the ($k-1$)-th cycle. Here, Pauli $Z$ measurements are executed on all qubits, which yield the measurement result vector $\bm{z}_{k-1}=[\langle Z_{1}\rangle_{k-1},\langle Z_{2}\rangle_{k-1},\dots,\langle Z_{N}\rangle_{k-1}]^{\top}$ (where $Z_{i}$ denotes the Pauli $Z$ for the qubit $i$). Returning to step (ii) of the $k$-th cycle, the vector $\bm{z}_{k-1}$ is provided back into the reservoir through the feedback connection. The operator $\mathcal{R}_{i,j}(a_{\mathrm{fb}}z_{k-1}^{\alpha})$ is employed for feedback with a weight $a_{\mathrm{fb}}$, which encodes all components of $\bm{z}_{k-1}$ according to the configuration shown in Fig. 2(a). Although the projective measurement strongly perturbs the quantum state, the feedback connections operate to systematically restore the memory of the provided input history.
@@
Subsequent to the input and feedback processes, a fixed eight-qubit gate $U_{\mathrm{res}}$ is applied in step (iii) to generate entanglement among all qubits. This gate serves as the reservoir component of our QRC architecture, which simultaneously transforms the input and feedback information in a non-linear manner. We employ a Haar random unitary matrix for $U_{\mathrm{res}}$ in the following calculations: refer to Appendix A for a hardware-efficient implementation of $U_{\mathrm{res}}$. Finally, step (iv) concludes the $k$-th cycle with projective measurements on all qubits under the assumption of negligible statistical fluctuations. The measurement result $\bm{z}_{k}$ is further cycled back to the reservoir via the feedback gate for processing the subsequent $(k+1)$-th input. We note that additional feedback layers can be integrated to capture long-term correlations in data, as illustrated in Appendix B.
@@
Figure 2: (a) Circuit diagram of the feedback-driven QRC architecture at the $k$-th cycle. The red and blue background colors represent the gates for the input [step (i)] and feedback [step (ii)] parts, respectively. The input value $s_{k}$ is provided to qubits $1$ and $2$ via $\mathcal{R}_{1,2}(a_{\mathrm{in}}s_{k})$ with the input weight $a_{\mathrm{in}}$. The feedback of the measurement result $\bm{z}_{k-1}$ is given to qubits $3$ to $8$ via $\mathcal{R}_{i,j}(a_{\mathrm{fb}}z_{k-1}^{\alpha})$. Subsequently, the entanglement among all qubits is introduced by the reservoir gate $U_{\mathrm{res}}$ [step (iii)], and all qubits are then measured in the projective manner at the end of the circuit [step (iv)]. (b) Circuit representation of the two-qubit operator $\mathcal{R}$ defined in Eq. (1).
@@
It is worth highlighting a key advantage of our architecture in terms of time complexity, which arises from the intrinsic independence of quantum circuits over different cycles. In our QRC, the quantum state at the $k$-th cycle depends exclusively on the current input value $s_{k}$ and the preceding measurement outcome $\bm{z}_{k-1}$, explicitly represented as $|\psi(s_{k},\bm{z}_{k-1})\rangle$. Even after the measurement, this state can be efficiently reconstructed by applying a limited set of gates, leveraging $\bm{z}_{k-1}$ stored in external memory; there is no need to re-execute the $(k-1)$-th and prior cycles when addressing the $k$-th input. Consequently, our protocol offers remarkable operational flexibility, enabling features such as starting from an intermediate cycle or running different cycles on different devices. In addition, as the circuit terminates with measurements after each individual cycle, a coherence time sufficient for the execution of just one cycle is adequate for our protocol. This aspect stands in stark contrast to other implementations depicted in Figs. 1(a) and 1(c), which, in order to process temporal data up to the $L$-th input cycle, necessitate $L$ repetitions of the reservoir operations to construct the history-dependent quantum state $|\psi(s_{k},s_{k-1},s_{k-2}\dots)\rangle$. Assuming a single cycle duration of $T$, a limitation on the input data length $L$ is thus imposed to ensure that the operation time $LT$ remains within the bound of the coherence time. In the protocol shown in Fig. 1(b), the required coherence time is reduced to $\tau T$ by limiting the number of input repetitions at each cycle to $\tau$; however, the need for a large $\tau$ to enhance performance is still in conflict with the temporal constraint imposed by the coherence time. In light of this comparison, our feedback-driven protocol necessitates the most minimal coherence time among these approaches, making it particularly well-suited for hardware implementations.
@@ II.2 Learning procedure
In the QRC protocol, the final output $y_{k}$ at the $k$-th cycle is obtained through a linear transformation of the measurement result $\bm{z}_{k}$. Here, a specific task for the QRC is formulated by setting a target output $\{\bar{y}_{k}\}$ based on the input sequence $\{s_{k}\}$. Our objective in the output part is to construct a regression model from the collection of measurement outcomes that accurately approximates the target output over various cycles. As a prerequisite, we define $\bm{z}_{0}$, required for the first cycle ($k=1$), as a random vector with each element sampled uniformly from the interval $[0,1]$. The influence of this initial condition is disregarded by discarding the initial $l_{\mathrm{w}}$ cycles.
@@
Following the training phase, we construct an $[l_{\mathrm{ts}}\times(N+1)]$-dimensional matrix ${X}_{\mathrm{ts}}$ for testing in a manner analogous to Eq. (2), using the measurement outcomes from the ($l_{\mathrm{w}}+l_{\mathrm{tr}}+1$)-th cycle to the ($l_{\mathrm{w}}+l_{\mathrm{tr}}+l_{\mathrm{ts}}$)-th cycle. The output is then calculated as $\bm{y}_{\mathrm{ts}}={X}_{\mathrm{ts}}\bm{w}$ using the trained weight. The similarity between the output $\bm{y}_{\mathrm{ts}}$ and the target output $\bar{\bm{y}}_{\mathrm{ts}}$ in the testing phase signifies the reservoir performance. As a quantitative metric, we utilize the determination coefficient
$R^{2}=\frac{\mathrm{cov}^{2}(\bar{\bm{y}}_{\mathrm{ts}},\bm{y}_{\mathrm{ts}})}{\sigma^{2}(\bar{\bm{y}}_{\mathrm{ts}})\sigma^{2}(\bm{y}_{\mathrm{ts}})},$ (4)